evaluate ∫ sqrt(x){x^3 – 2/x} dx

evaluate
∫ sqrt(x){x^3 – 2/x} dx

Grade:12

1 Answers

Jitender Singh IIT Delhi
9 years ago
Ans:
Hello Student,
Please find answer to your question below

I = \int \sqrt{x}(x^3-\frac{2}{x}) dx
I = \int (x^{7/2}-\frac{2}{\sqrt{x}}) dx
I = \int x^{7/2}dx-\int \frac{2}{\sqrt{x}} dx
I = \frac{x^{\frac{7}{2}+1}}{\frac{7}{2}+1}- \int \frac{2}{\sqrt{x}}dx
I = \frac{2x^{\frac{9}{2}}}{9} - \int \frac{2}{\sqrt{x}}dx
I = \frac{2x^{\frac{9}{2}}}{9} - 4\sqrt{x}+c